OpenStudy (anonymous):

f(X)=x^3-x^2-4x+4...Factor this polynomial completely over the set of complex numbers

OpenStudy (anonymous):

always check first if \(f(1)=0\)

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

replace \(x\) by \(1\)

OpenStudy (anonymous):

do it in your head \[f(1)=1-1-4+4=0\]

OpenStudy (anonymous):

yes its equal to zero

OpenStudy (anonymous):

once you see that \(f(1)=0\) you can factor as \[x^3-x^2-4x+4.=(x-1)(something)\]

OpenStudy (anonymous):

you can find the "something" by division or synthetic division or thinking the "something" will be a quadratic, and you can find the zero of that using the quadratic formula if you have to

OpenStudy (anonymous):

ok i'll try the quadratic formula so....is a=-1 b=-4 c=4 ?

OpenStudy (anonymous):

did you factor yet?

OpenStudy (anonymous):

you have to factor before you can use the quadratic formula

OpenStudy (anonymous):

factor what?

OpenStudy (anonymous):

\[x^3-x^2-4x+4=(x-1)(something)\] you need the "something"

OpenStudy (anonymous):

you have to divide \(x^3-x^2-4x+4\) by \(x-1\) to find it

OpenStudy (luigi0210):

You could also factor by grouping, if I'm not mistaken.

OpenStudy (anonymous):

ok so i used synthetic division and i got (x-1)(2x^2-4x+4), correct?

OpenStudy (anonymous):

where did the 2 come from?

OpenStudy (anonymous):

the leading coefficient is 1, there is no 2 in it

OpenStudy (anonymous):

yea but dont in synthetic division you multiply -1 by 1 and than on the top i have -1 + -1=-2

OpenStudy (anonymous):

1 -1 -4 4 1 _____________ 1

OpenStudy (anonymous):

leading coefficient is 1

OpenStudy (anonymous):

ok i see what you mean i put -1 as the leading coefficient

OpenStudy (anonymous):

no you put 2

OpenStudy (anonymous):

ok well....now i got 1 0 -4 0

OpenStudy (anonymous):

1 -1 -4 4 1 -1 0 -4 ____________ 1 0 -4 0

OpenStudy (anonymous):

so it is \[(x-1)(x^2-4)\]

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

yw